OpenStudy (anonymous):

Probe that: tan A - cotan A = -2 cotan 2A

OpenStudy (anonymous):

I didn't understand your reply...

OpenStudy (perl):

now we know that cos(2A) = cos^2(A) - sin^2(A)

OpenStudy (perl):

and we know that sin(2A) = 2 sin A cos A , these are trig identities

OpenStudy (perl):

tan A - cot A = sin A / cos A - cos A / sin A = ( sin^2 (A) - cos^2(A) ) / ( cos (A) sin(A) ) = - ( cos^2(A) - sin^2(A)) / (cos(A) sin (A)) = - ( cos(2A) ) / ( cos A sin A ) = - 2( cos (2A) ) / ( 2 cos A sin A ) = -2cos(2A) / ( sin (2A) = -2 cot (2A)

OpenStudy (anonymous):

OK, perl, now I understood your explanation. Thank you very much! Bye!